A car moving in a straight line starts at X=0 at t=0. It passesthe point x=25.0 m with a speed of 11.0 m/s at t=3.0 s. It passesthe point x=385 with a speed of 45.0 m/s at t=20.0 s. Find theaverage velocity and the average accelerationbetween t=3.0 s and20.0 s.
Answer #1

x1 = 0 x2 = 25.0 m

t1 = 0 x2= 3.00 sec at speed 11.0 m/s

passes through point x = 385.0 m at speed 45.0 m/s

t = 20.0 sec

A. average velocity = (385.0 m - 25.0 m) / (20 sec - 3 sec) = 360.0 m / 17 sec = 21.2 m/s


B. average acceleration = (45 m/s - 11 m/s) / (20 sec - 3 sec) = 34 m/s / 2 sec = 2 m/s^2